OpenStudy (narissa):

please help The length (L) of a photograph is 8mm longer than its width (W). Its perimeter (P) is more than 60mm. 𝐏 = 𝟐𝐖 + 𝟐𝐋 a) Write an equation for the length in terms of the width L = _____________________ b) Use these expressions to write an inequality based on the given information. c) Solve the inequality clearly indicating the length and width of the photograph. d) Write your answer in a complete sentence

OpenStudy (narissa):

@campbell_st

OpenStudy (narissa):

@freckles

OpenStudy (narissa):

@DanJS

OpenStudy (narissa):

@Loser66

OpenStudy (loser66):

Replace L = 8 + W into P.

OpenStudy (narissa):

im confused so p=2w+2(8)

OpenStudy (narissa):

@Loser66

OpenStudy (narissa):

@Mehek14

OpenStudy (loser66):

I said L = 8+W P = 2W + 2L = 2W +2(8+W) =??

OpenStudy (narissa):

16?

OpenStudy (narissa):

@jim_thompson5910

OpenStudy (narissa):

@YoungStudier

OpenStudy (narissa):

@SithsAndGiggles